# HRFT: Mining High-Frequency Risk Factor Collections End-to-End via Transformer

### v1

@@ Introduction
EVALUATION
@@
Experimental Settings
Datasets
Our experiments are conducted on HFT data from the U.S. and China stock market (see Table 1). The stock data was obtained from wind [^1] https://www.wind.com.cn/. Specifically, we analyze the $W$-dimensional trading data $X\in\mathbb{R}^{W}$ of the 1-min trading level of constituent stocks of the HS300 index and the S&P500 index. We preprocess raw stock open/close/high/low/volume/vwap features using word expressions. It’s important to highlight that our method employs an ascending variable sampling approach for features. This implies that instances such as $x_{2}+x_{3}$ will not occur, but rather samples like $x_{1}+x_{2}+x_{3}$ will be generated. However, we have the flexibility to set the weight in front of $x_{1}$ to zero. The target value is the one-day-ahead RV, defined as $RV(t,j;n)=\sum_{j=1}^{n}(\ln P_{t,j}-\ln P_{t,j-1})^{2}$, where $n$ is the number of intraday trading intervals and $P_{i,j}$ is the closing price of the $j$-th interval of the $i$-th trading day (Andersen and Bollerslev 1998).
Table 1: Market Data Overview
@@
Compared Approaches
We compare IRFT with advanced methods from SRBench(La Cava et al. 2021) in generating HF risk factor expressions.
  • DL-based methods: DSR(Petersen et al. 2019) employs a parameterized neural network to model the distribution of a mathematical expression, represented as a sequence of operators and variables in the corresponding expression tree.
@@
  • ML-based methods: To evaluate the model more comprehensively, we compare IRFT with the end-to-end ML methods in SRBench. These methods take a week’s HF stock features as input and predict one-day-ahead RV. AdaBoost ensembles decision trees to directly predict stock trends. Lasso Lars is a linear regression method that combines Lasso and LARS. Random Forest is an ensemble learning method that averages the regression results from multiple decision trees. The other ML methods are SGD Regressor (Bottou 2010), MLP and Linear Regression. These ML models do not generate formulaic HF risk factor expressions, but only output volatility prediction values. The hyperparameters are set by SRBench.
Evaluation Metrics
@@
$R^{2}=1-\frac{\sum_{i=1}^{M_{\text{test}}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{M_{\text{test}}}(y_{i}-\bar{y})^{2}}$ (1)
$\boldsymbol{IC^{*}.}$
It is the cross-section correlation coefficient of the HF risk factor values and the one-day-ahead RV values of trading stocks, which ranges from $[-1,1]$.
$IC^{*}=\bar{\sigma}\left(\sum_{i=1}^{k}w_{i}f_{i}(X),y_{i}\right)$ (2)
@@
$RankIC^{*}=\bar{\sigma}\left(r\left(\sum_{i=1}^{k}w_{i}f_{i}(X)\right),r(y_{i})\right)$ (3)
@@
$IR^{*}=\frac{\bar{Ret}_{t}^{IC}}{\text{std}(Ret_{t}^{IC})}$ (4)
Figure 3: The Pareto plot compares our framework with baselines like DL-based methods, ML-based methods, and GP-based methods from SRBench about the average inference time and test performance.
Table 2: Main results of S&P500 Index and HS300 Index. Values after ”$\pm$” represents the standard deviation, and the rest are the mean values, while the rest are mean values. “↑” indicates that higher values are better.
Method S&P 500 HS 300
$IC^{*}$ (↑) Rank $IC^{*}$ (↑) $IR^{*}$ (↑) $IC^{*}$ (↑) Rank $IC^{*}$ (↑) $IR^{*}$ (↑)
DSR $0.0437\pm 0.0054$ $0.0453\pm 0.0071$ $0.2506\pm 0.0246$ $0.0427\pm 0.0088$ $0.0456\pm 0.0059$ $0.3395\pm 0.0637$
GPLEARN $0.0250\pm 0.0099$ $0.0547\pm 0.0073$ $0.4876\pm 0.0307$ $0.0494\pm 0.0062$ $0.0480\pm 0.0063$ $0.3600\pm 0.0368$
GENEPRO $0.0470\pm 0.0068$ $0.0550\pm 0.0163$ $0.2098\pm 0.0339$ $0.0444\pm 0.0049$ $0.0528\pm 0.0090$ $0.4787\pm 0.0453$
Table 3: Top 5 HF risk factor expressions based on $IC^{*}$ values in the HF risk factor collections (S&P500 Index).
No. HF risk factor $IC^{*}$
1 $0.6*\log((2709.5-0.9\cdot\sin(0.4\cdot\text{close}-2.0)+16.7-40.1/((0.2\cdot\text{low}+0.4))))-0.4)+0.2$ 0.0820
2 $-8.4+8.2/(-0.7*(-0.1-0.1/(1.9-0.5\cdot\text{close})))$ 0.0641
3 $1.02.8*\sqrt{0.1/((-0.2\cdot\text{open}-11.7*(13.7-2.6\cdot\text{open})+1.0)^{2})}-$ 1.0 0.0635
4 $1.3$ ( 0.2 ∗ | - 0.1 ⋅ low + 5.1 | + 0.2 ) + 1.3 0.0558
5 $0.8/((-0.1\cdot\text{open}-(0.1\cdot\text{high}+3.6)*(-3.0*(5.2-1.5\cdot\text{volume})*(40.0-7.0\cdot\text{low})+2046.9)))-0.5$ 0.0522
Figure 4: (a) Performance of different methods on S&P500 Index (U.S. market). (b) A comparison of different factor generation methods using $IC^{*}$, $RankIC^{*}$ and $IR^{*}$ for various factor pool sizes.
Figure 5: (a) The correlation of the HF risk factor collections on S&P500 Index. (b) Portfolio simulations on the S&P500 Index: a backtesting comparison.
Comparison across all HF risk factor generators (RQ1)
We evaluate the performance of three baseline methods on HFT data sets derived from the constituents of the S&P500 index. The results, in terms of inference time, expression complexity, and prediction performance ($R^{2}$), are presented using box plots in Fig. 4.
@@
Comparison of formulaic generators with varying pool capacity (RQ2)
@@
Investment Simulation (RQ3)
During backtesting period, we analyze factor set heat maps (Figures 5(a) and 7(a)) and select 10 independent risk factors(the more information they contain). We calculate risk factor scores for all assets in the portfolio ($k=30$ stocks in our paper) for each adjustment period (each trading day). The factor score for the $s$th stock is $a(s)=\sum_{n=1}^{N}w(n)V(s,n)$, where $w(n)$ is the weight of the $n$th factor ($\text{IC}^{*}$ value) and $V(s,n)$ is the $n$th factor value of the $s$th stock. These factors help select high-yielding stocks for the portfolio. As shown in Figure 5(b), we record cumulative net returns over a one-year test period. Despite the fact that our framework does not explicitly optimize for absolute return, it still performs well in the backtest, even during the test period when the international environment is unstable and the domestic economy slows down. Compared with other methods, our framework can still achieve the highest profit, exceeding the second-place method GENEPRO by 21%. Table 3 presents the top five HF risk factors selected from the HF risk factor collections, based on their $IC^{*}$ values. These factors, which include arbitrary constant terms, are highly flexible and can be adapted to any range of values. For simplicity, we round the constant terms to one decimal place.
Conclusion
@@
Table 4: List of all operators employed within our framework.
@@
Table 5: Detailed parameter settings of IRFT.
Parameter Description Value
W=5(S&P500) W=5(SSE50)
W_max Max input dimension 10
b_max Max binary operators 5+W 5+W
O_b Binary operators add,sub,mul,div
u_max Max unary operators 5 5
O_u Unary operators inv,abs,sqr,sqrt,sin,cos,tan,atan,log,exp
M_min Min number of points 10D
M_max Max number of points 200
c_max Max number of candidate function clusters 10
Figure 6: (a) Performance of different methods on HS300 Index(the China market). (b) A comparison of different factor generation methods using $IC^{*}$ , $RankIC^{*}$ and $IR^{*}$ for various factor pool sizes.
Figure 7: (a) The correlation of the HF risk factor collections on HS300 Index. (b) Portfolio simulations on the HS 300 Index: a backtesting comparison.
Collection of operators
@@
The detailed parameter set description
Performance of different methods under the China HFT market
As depicted in Fig. 6, our method demonstrates a significant advantage in terms of inference time, outperforming the slowest method, DSR six times. The concentrated distribution of inference times indicates high stability and efficiency of our method. GPLEARN, while faster than DSR, exhibit more dispersed distributions, suggesting their performance is more susceptible to variations in the stock data set. In terms of expression complexity, DSR generates the most concise expressions, suggesting its ability to find simple and effective HF risk factor expressions. Regarding prediction performance, our method generates HF risk factor expressions with a value of $R^{2}$ of 0.87, significantly outperforming other methods. The next best method is GPLEARN, followed by GENEPRO, with DSR performing the worst.
Comparison of formulaic generators with varying pool capacity(HS300 Index).
We further extend our analysis to evaluate the performance of different methods in generating HF risk factor collections for the China stock market. As depicted in Fig. 6, our method, IRFT, consistently exhibits the highest values of $IC^{*}$, $RankIC^{*}$, and $IR^{*}$ across all factor pool sizes. This indicates its superior predictive ability in generating HF risk factors. The performance of the other three methods does not differ significantly.
HF risk collections for investment simulation(SSE50 Index).
As depicted in Fig. 7(a), we have implemented a filtering process to delete HF risk factors with high correlation. Consequently, the HF risk factor collection retains 10 unique HF risk factors. As illustrated in Fig. 7(b), we recorded the cumulative net value of all strategies over a one-year test period in the China HFT market. Our framework outperforms other methods, achieving the highest profit and exceeding other SR methods by 30%. Next, GPLEARN obtains higher portfolio returns than the remaining two benchmarks. The one that receive the lowest return is DSR. The backtest results demonstrate that our method still has maintained stable growth throughout the backtest period.

### v5

@@ 3.3.4. Inference time
Figure 3. The Pareto plot compares our framework with baselines like DL-based methods, ML-based methods, and GP-based methods from SRBench about the average inference time and test performance.
Table 2. Main results of S&P500 Index and HS300 Index. Values after $\pm$” represents the standard deviation, and the rest are the mean values, while the rest are mean values. “↑” indicates that higher values are better.
Method S&P 500 HS 300
$IC^{*}$ (↑) Rank $IC^{*}$ (↑) $IR^{*}$ (↑) $IC^{*}$ (↑) Rank $IC^{*}$ (↑) $IR^{*}$ (↑)
DSR $0437\pm 0.0054 0.0$$0453\pm 0.0071$ $2506\pm 0.0246 0.0$$427\pm 0.0088 0.0$$0456\pm 0.0059$ $0.3395\pm 0.0637$
GPLEARN $0250\pm 0.0099 0.0$$0547\pm 0.0073$ $4876\pm 0.0307 0.0$$494\pm 0.0062 0.0$$0480\pm 0.0063$ $0.3600\pm 0.0368$
GENEPRO $0470\pm 0.0068 0.0$$0550\pm 0.0163$ $2098\pm 0.0339 0.0$$444\pm 0.0049 0.0$$0528\pm 0.0090$ $0.4787\pm 0.0453$
4. EVALUATION
@@
Table 3. Top 5 HF risk factor expressions based on $IC^{*}$ values in the HF risk factor collections (S&P500 Index).
No. HF risk factor $IC^{*}$
1 $6*\log((2709.5-0.9\cdot\sin(0.4\cdot\text{close}-2.0)+16.7-40.1/((0.2\cdot\text{low}+0.4))))-0.4)+0.2 0.0$820
2 $8.4+8.2/(-0.7*(-0.1-0.1/(1.9-0.5\cdot\text{close})))$ 0.0641
3 $2.8*\sqrt{0.1/((-0.2\cdot\text{open}-11.7*(13.7-2.6\cdot\text{open})+1.0)^{2})}-1.0$ 0.0635
4 $1.3*(0.2*\lvert-0.1\cdot\text{low}+5.1\rvert+0.2)+1.3$ 0.0558
5 $8/((-0.1\cdot\text{open}-(0.1\cdot\text{high}+3.6)*(-3.0*(5.2-1.5\cdot\text{volume})*(40.0-7.0\cdot\text{low})+2046.9)))-0.5 0.0$522
Figure 4. (a) Performance of different methods on S&P500 Index (U.S. market). (b) A comparison of different factor generation methods using $IC^{*}$, $RankIC^{*}$ and $IR^{*}$ for various factor pool sizes.
4.1. Experimental Settings
4.1.1. Datasets
Our experiments are conducted on HFT data from the U.S. and China stock market (see Table 4). The stock data was obtained from wind [^1] https://www.wind.com.cn/. Specifically, we analyze the $W$-dimensional trading data $X\in\mathbb{R}^{W}$ of the 1-min trading level of constituent stocks of the HS300 index and the S&P500 index. We preprocess raw stock open/close/high/low/volume/vwap features using word expressions. It’s important to highlight that our method employs an ascending variable sampling approach for features. This implies that instances such as $x_{2}+x_{3}$ will not occur, but rather samples like $x_{1}+x_{2}+x_{3}$ will be generated. However, we have the flexibility to set the weight in front of $x_{1}$ to zero. The target value is the one-day-ahead RV, defined as $RV(t,j;n)=\sum_{j=1}^{n}(\ln P_{t,j}-\ln P_{t,j-1})^{2}$, where $n$ is the number of intraday trading intervals and $P_{i,j}$ is the closing price of the $j$-th interval of the $i$-th trading day (Andersen and Bollerslev, 1998).
Table 4. Market Data Overview
@@ 4.1.1. Datasets
4.1.2. Compared Approaches
We compare IRFT with advanced methods from SRBench(La Cava et al., 2021) in generating HF risk factor expressions.
  • DL-based methods: DSR(Petersen et al., 2019) employs a parameterized neural network to model the distribution of a mathematical expression, represented as a sequence of operators and variables in the corresponding expression tree.
@@ 4.1.2. Compared Approaches
  • ML-based methods: To evaluate the model more comprehensively, we compare IRFT with the end-to-end ML methods in SRBench. These methods take a week’s HF stock features as input and predict one-day-ahead RV. AdaBoost ensembles decision trees to directly predict stock trends. Lasso Lars is a linear regression method that combines Lasso and LARS. Random Forest is an ensemble learning method that averages the regression results from multiple decision trees. The other ML methods are SGD Regressor (Bottou, 2010), MLP and Linear Regression. These ML models do not generate formulaic HF risk factor expressions, but only output volatility prediction values. The hyperparameters are set by SRBench.
4.1.3. Evaluation Metrics
@@
(1) $R^{2}=1-\frac{\sum_{i=1}^{M_{\text{test}}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{M_{\text{test}}}(y_{i}-\bar{y})^{2}}$
$\boldsymbol{IC^{*}}$
It is the cross-section correlation coefficient of the HF risk factor values and the one-day-ahead RV values of trading stocks, which ranges from [1,1][-1,1].
(2) $IC^{*}=\bar{\sigma}\left(\sum_{i=1}^{k}w_{i}f_{i}(X),y_{i}\right)$
@@
(3) $RankIC^{*}=\bar{\sigma}\left(r\left(\sum_{i=1}^{k}w_{i}f_{i}(X)\right),r(y_{i})\right)$
@@
(4) $IR^{*}=\frac{\bar{Ret}_{t}^{IC}}{\text{std}(Ret_{t}^{IC})}$
Figure 5. (a) The correlation of the HF risk factor collections on S&P500 Index. (b) Portfolio simulations on the S&P500 Index: a backtesting comparison.
4.2. Comparison across all HF risk factor generators (RQ1)
We evaluate the performance of three baseline methods on HFT data sets derived from the constituents of the S&P500 index. The results, in terms of inference time, expression complexity, and prediction performance ($R^{2}$), are presented using box plots in Fig. 4(a).
@@ 4.2. Comparison across all HF risk factor generators (RQ1)
4.3. Comparison of formulaic generators with varying pool capacity (RQ2)
@@
4.4. Investment Simulation (RQ3)
During backtesting period, we analyze factor set heat maps (Fig. 5(a) and select 10 independent risk factors(the more information they contain). We calculate risk factor scores for all assets in the portfolio ($k=30$ stocks in our paper) for each adjustment period (each trading day). The factor score for the $s$th stock is $a(s)=\sum_{n=1}^{N}w(n)V(s,n)$, where $w(n)$ is the weight of the $n$th factor ($\text{IC}^{*}$ value) and $V(s,n)$ is the $n$th factor value of the $s$th stock. These factors help select high-yielding stocks for the portfolio. As shown in Fig. 5(b), we record cumulative net returns over a one-year test period. Despite the fact that our framework does not explicitly optimize for absolute return, it still performs well in the backtest, even during the test period when the international environment is unstable and the domestic economy slows down. Compared with other methods, our framework can still achieve the highest profit, exceeding the second-place method GENEPRO by 21%. Table 3 presents the top five HF risk factors selected from the HF risk factor collections, based on their $IC^{*}$ values. These factors, which include arbitrary constant terms, are highly flexible and can be adapted to any range of values. For simplicity, we round the constant terms to one decimal place.
5. Related Work
Symbolic Regression. Symbolic Regression (SR) focuses on identifying mathematical expressions that best describe relationships (Kamienny et al., 2022). Mining quantitative factors is essentially an SR task, extracting accurate mathematical patterns from historical trading data. The dominant approach is Genetic Programming (GP) (Cui et al., 2021), but it has two drawbacks: slow execution (Langdon, 2011), due to the expansive search space requiring numerous generations to converge, and a tendency towards overfitting (Fitzgerald et al., 2013), dependent on input fitness. Given the time-sensitive nature of quantitative trading, we aim to generate risk factors that effectively capture short-term market volatility.
Language Models. The pre-trained Finbert(Liu et al., 2021) or other pre-trained Fingpt(Yang et al., 2023; Liu et al., 2023; Xu et al., 2025c, b) mainly focus on semantic processing. However, the task of generating formulaic factors primarily focuses on tackling numbers and symbols, rather than comprehending extensive vocabularies. They encounter difficulties in complex reasoning tasks, like predicting formulas for large numbers(Mishra et al., 2022). In pretraining, models struggle with rarely or never seen symbols. And they can only deal with sequences. We treat HF risk factors as a language for explaining market volatility trends and train a Transformer. This model offers two benefits: They are fast because of leveraging previous experience with inference being a single forward pass. And they are less over fitting-prone, as they do not require loss optimization based on inputs.
6. Conclusion
@@
Acknowledgements.
This research was supported by the Beijing Natural Science Foundation (Grant No. 4232039), the Youth Fund of the Computer Network Information Center, Chinese Academy of Sciences (Grant No. 24YF07), and the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500103).
Appendix A APPENDIX
Table 5. List of all operators employed within our framework.
@@ Appendix A APPENDIX
Figure 6. (a) Performance of different methods on HS300 Index (China market). (b) A comparison of different factor generation methods using $IC^{*}$, $RankIC^{*}$ and $IR^{*}$ for various factor pool sizes on HS300 Index.
Figure 7. (a) The correlation of the HF risk factor collections on HS300 Index. (b) Portfolio simulations on the HS300 Index: a backtesting comparison.
A.1. Collection of operators
@@
A.2. Performance of different methods under the China HFT market
As depicted in Fig. 6(a), our method demonstrates a significant advantage in terms of inference time, outperforming the slowest method, DSR six times. The concentrated distribution of inference times indicates high stability and efficiency of our method. GPLEARN, while faster than DSR, exhibit more dispersed distributions, suggesting their performance is more susceptible to variations in the stock data set. In terms of expression complexity, DSR generates the most concise expressions, suggesting its ability to find simple and effective HF risk factor expressions. Regarding prediction performance, our method generates HF risk factor expressions with a value of $R^{2}$ of 0.87, significantly outperforming other methods. The next best method is GPLEARN, followed by GENEPRO, with DSR performing the worst.
A.3. Comparison of formulaic generators with varying pool capacity(HS300 Index).
We further extend our analysis to evaluate the performance of different methods in generating HF risk factor collections for the China stock market. As depicted in Fig. 6(b), our method, IRFT, consistently exhibits the highest values of $IC^{*}$, $RankIC^{*}$, and $IR^{*}$ across all factor pool sizes. This indicates its superior predictive ability in generating HF risk factors. The performance of the other three methods does not differ significantly.
A.4. HF risk collections for investment simulation(SSE50 Index).
As depicted in Figure 7(a), we have implemented a filtering process to delete HF risk factors with high correlation. Consequently, the HF risk factor collection retains 10 unique HF risk factors. As illustrated in Figure 7(b), we recorded the cumulative net value of all strategies over a one-year test period in the China HFT market. Our framework outperforms other methods, achieving the highest profit and exceeding other SR methods by 30%. Next, GPLEARN obtains higher portfolio returns than the remaining two benchmarks. The one that receive the lowest return is DSR. The backtest results demonstrate that our method still has maintained stable growth throughout the backtest period.